# Bidirectional Decoding: Improving Action Chunking via Closed-Loop Resampling

### v2

@@ 5.1 One-Dimensional Diagnostic Experiments
Result. As shown in Fig. 5, when the environment is deterministic ($\delta=0$), larger action horizons capture the expert distribution better, consistent with Corollary 2. With an action horizon of 10, the learner achieves zero total variation distance with the expert distribution. Conversely, when the environment is highly stochastic $\delta=0.8$, an action horizon of 1 outperforms all other learners, corroborating with Corollary 3. With moderate noise $\delta=0.4$, there is no discernible monotonic pattern due to the tradeoff revealed in Proposition 1. Refer to Appendix A for more detailed results.
@@ 5.2.1 Comparison with Existing Inference Methods
  • Exponential Moving Average (EMA) (Zhao et al., 2023): Smooth action chunking by averaging a new prediction $a$ with the previous one $\hat{a}$ for each overlapping step $a_{t}=\lambda a_{t}+(1-\lambda){\hat{a}}_{t}$. This method is also known as temporal ensembling. By default, we set $\lambda=0.5$.
@@ 5.2.3 Generality and Efficiency of BID
Stochastic Noise 0.0 1.0 1.5
Vanilla Open-Loop 61.0 39.0 19.4
BID Open-Loop 65.2 39.8 21.4
Vanilla Closed-Loop 52.0 50.4 44.2
BID Closed-Loop 56.6 54.8 54.4
Table 1: Success rates of VQ-BeT on the Push-T task under various conditions. BID consistently outperforms the vanilla counterpart. Closed-loop BID is particularly advantageous in stochastic settings.
Sample Size Success (%) Time (ms)
1  (vanilla) 52.0 12.6
8  (ours) 53.8 25.6
16  (ours) 56.6 26.4
32  (ours) 56.6 27.3
@@ 6 Conclusion
Limitations. One major limitation of BID lies in its computational complexity. While the decoding process can be parallelized on modern GPUs, it may remain prohibitive for high-frequency operations on low-cost robots. Designing algorithms that can generate quality yet diverse action chunks under batch size constraints can be an interesting avenue for future research. Additionally, our analysis and method have been limited to policies with short context lengths, driven by their empirical effectiveness with limited human demonstrations. Developing techniques capable of learning robust long-context policies can be another compelling direction for future research.
@@
To facilitate the reproducibility of our results, we release our code for the experiments with Diffusion Policy at https://github.com/YuejiangLIU/bid_diffusion and our code for the experiments with VQ-BET at https://github.com/Jubayer-Hamid/bid_lerobot. Additionally, we provide our detailed experimental setups in Appendix B, and complete proofs of our theoretical analysis in Appendix C.
@@ A.1 One-dimensional Simulations
Figure 11: Effect of prediction horizon ($h$) and context length ($c$) on diffusion policies in the Push-T task. The baseline is set at $h=2$ and $c=2$, with $\Delta h=h-c=0$. Extending the prediction horizon ($h>2$) consistently improves performance, whereas extending the context length ($c>2$) can cause substantial performance declines. Each model is trained for $5k$ epochs. Results are averaged over the last five checkpoints.
@@
Setup. Our work builds on the premise that the action horizon is longer than the context length, as commonly designed for recent policies. While BID mitigates the inherent limitations of this design choice through test-time decoding, an important question remains: could extending the context length yield stronger policies? To understand this, we trained diffusion policies with varying combinations of action horizons and context lengths on the Push-T task. Specifically, we use a short context length ($c=2$) and a short action horizon ($h=2$) as our baseline, and incrementally increase these parameters to larger values ${6,10,14}$ to assess their impact.
@@ A.3 Ablation Study of Forward Contrast
Figure 13: Effect of the decay rate for the exponential moving average. In each task, we measure the relative performance among different decay rates. The optimal decay rate varies by task, leading to a practical challenge of identifying a universal temporal ensembling strategy (Zhao et al., 2023).
@@ B.1.1 Environment Details
Our implementation of BID for VQ-BeT (Lee et al., 2024) is built upon the code of LeRobot (Cadene et al., 2024). We use the best public checkpoint as the strong policy and a checkpoint trained for 100k iterations as the weak policy. Since BID requires sample diversity, we set the temperature as 0.5 for our methods, while vanilla sampling uses the default temperature of 0.1.
@@ B.2.1 Dynamic Placing
Following previous works (Chi et al., 2023; Prasad et al., 2024), we use a Franka Panda as the robot hardware and the vision-based diffusion policy for its operation. The robot is equipped with two cameras: one ego-centric camera mounted at the wrist of the robot, one third-person camera mounted at a static bracket. Both cameras provide visual observations at a resolution of $256\times 256$ pixels. The robot operates at a frequency of 10 Hz, with a prediction horizon of 16 time steps.
@@ Results.
Appendix C Proofs
First, we establish the following lemma which will help us compare different function classes:
@@ Appendix C Proofs
where we used $c^{*}(X):=\arg\min_{c}\mathbb{E}_{X}[\mathcal{L}(f(X,Y),c)|X]$. We get the inequality by recognizing that $\mathbb{R}\subsetneq G$. On the other hand, the left hand side is greater than or equal to the right hand side. For any $g(X)$, we have:
@@
C.1 Definitions
Note that the informal definitions provided in Sec. 3.2 for expected observation advantage and maximum inference disadvantage have some deviations from the mathematical definitions. The informal definitions only attempt to provide intuition for what these terms might mean but they are not sufficient to describe the full construction.
To define the terms formally, we, first, analyze the effect of reducing context horizon. We show that, provided action horizon is constant, decreasing context horizon causes performance of the optimal policy to decrease.
Consider a $(c,h)$-policy whose probability of taking action $a_{t}$ at time $t$ in a chunk generated at $t$ is referred to as
@@ C.1 Definitions
We refer to the class of functions that accept $a_{t}$ and $s_{t-c-1:t}$ as inputs as $X_{+,+}$. Similarly, the class of functions that do not accept $a_{t}$ as inputs but accept $s_{t-c-1:t}$ as inputs is $X_{+}$. The function class that accepts only $s_{t-c:t}$ and not $s_{t-c-1}$ or $a_{t}$ as inputs are elements of $X_{-}$. Lastly, the function class that accepts $s_{t-c:t}$ and $a_{t}$ as inputs, but not $s_{t-c-1}$, are elements of $X_{-,-}$.
$\min_{\pi_{(c+1,h)}\in X_{+,+}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+1,h)},\pi^{*}\Big{|}C\right]$
$=\mathbb{E}_{a_{t}}\left[\min_{\pi_{(c+1,h)}^{\prime}\in X_{+}}\mathbb{E}_{s_{t-c-1}}\left[\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c+1,h)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t-c-1},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4)
$=\mathbb{E}_{a_{t}}\left[\mathbb{E}_{s_{t-c-1}}\left[\min_{\pi_{(c,h)}^{\prime}\in X_{-}}\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c,h)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t-c-1},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4)
$\leq\mathbb{E}_{a_{t}}\left[\min_{\pi_{(c,h)}^{\prime}\in X_{-}}\mathbb{E}_{s_{t-c-1}}\left[\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c,h)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t-c-1},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Jensen’s inequality)
$=\min_{\pi_{(c,h)}\in X_{-,-}}\mathbb{E}_{a_{t}}\left[\mathbb{E}_{s_{t-c-1}}\left[\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*})\Big{|}a_{t},s_{t-c-1},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4).
Use the law of total expectation to conclude. ∎
@@
$0\leq\alpha_{b}=\min_{\pi_{(c,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*}))\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (8)
where $C=\{s_{t-h-c:t-h},a_{t-h:t-1}\}$ and $G=\{a_{t},s_{t-k:t-h-c-1},s_{t-h+1:t},z_{t-k:t}\}$. In particular, $\alpha_{b}=0$ when $s_{t-h-d-c:t-h-c-1}$ can be deterministically inferred by $\pi_{(c,h)}$ or when the expert policy is independent of them. However, this is extremely unlikely since $\pi_{(c,h)}$ does not know the actions taken in those time steps (even more unlikely in a stochastic environment) and the expert’s action depends on the last $k$ time steps.
Definition (Maximum Inference Disadvantage ($\epsilon_{f}$)). Consider the maximum divergence that can be accumulated by the ${(c,h+d)}$-policy from not knowing the recent states at time steps $s_{t-h-d+1:t-h}$ and let that be $\epsilon_{f}$. More formally, we say that, for fixed $C$ from Proposition 1, any state in $\mathcal{S}^{-}$ and any $z_{t-k:t}$ and any $\hat{s}_{t-h-d+1:t-h}\neq{s}_{t-h-d+1:t-h}$:
@@
Intuitively, maximum inference disadvantage captures the maximum dependency on relative time steps whereas expected observation advantage captures the expected advantage from observing some given states.
@@
Proposition 7 (Forward Context is valuable).
Let $\mathcal{L}$ be a non-linear, convex function. Let $c<k$. Let $G:=\{a_{t},s_{t-k:t-c-1},s_{t},z_{t-k:t}\}$ and let $C:=\{s_{t-c:t-1},a_{t-1}\}$. Then,
@@
The proof is similar to that of Proposition 6. We refer to the class of functions that accept $a_{t}$ and $s_{t-c:t}$ as inputs as $X_{+,+}$. Similarly, the class of functions that do not accept $a_{t}$ as inputs but accept $s_{t-c:t}$ as inputs is $X_{+}$. The function class that accepts only $s_{t-c:t-1}$ and not $s_{t}$ or $a_{t}$ as inputs are elements of $X_{-}$. Lastly, the function class that accepts $s_{t-c:t-1}$ and $a_{t}$ as inputs, but not $s_{t}$, are elements of $X_{-,-}$.
$\min_{\pi_{(c,h)}\in X_{+,+}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*}\Big{|}C\right]$
$=\mathbb{E}_{a_{t}}\left[\min_{\pi_{(c,h)}^{\prime}\in X_{+}}\mathbb{E}_{s_{t}}\left[\mathbb{E}_{s_{t-k:t-c-1},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c,h)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4)
$=\mathbb{E}_{a_{t}}\left[\mathbb{E}_{s_{t}}\left[\min_{\pi_{(c-1,h+1)}^{\prime}\in X_{-}}\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c-1,h+1)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4)
$\leq\mathbb{E}_{a_{t}}\left[\min_{\pi_{(c-1,h+1)}^{\prime}\in X_{-}}\mathbb{E}_{s_{t}}\left[\mathbb{E}_{s_{t-k:t-c-1},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c-1,h+1)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Jensen’s inequality)
$=\min_{\pi_{(c-1,h+1)}\in X_{-,-}}\mathbb{E}_{a_{t}}\left[\mathbb{E}_{s_{t}}\left[\mathbb{E}_{s_{t-k:t-c-1},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c-1,h+1)},\pi^{*})\Big{|}a_{t},s_{t},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4).
Use the law of total expectation to conclude. ∎
@@
$0\leq\alpha_{f}=\min_{\pi_{(c,h+d)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h+d)},\pi^{*}))\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (10)
where $C=\{s_{t-h-d-c:t-h-d},a_{t-h-d:t-1}\}$ and $G=\{a_{t},s_{t-k:t-h-d-c-1},s_{t-h-d+1:t},z_{t-k:t}\}$. In particular, $\alpha_{f}=0$ when $\pi_{(c,h+d)}$ can infer $s_{t-h-d+1:t-h}$ perfectly i.e. when the environment is completely static with $P_{f}=1$. This makes sense–in the static environment, observing these states does not provide any advantage since the optimal $\pi_{(c,h+d)}$ can infer these states anyway using the actions taken at those time steps.
Note how this formal definition has some difference from the informal one. In particular, $\pi_{h}$ only observes $s_{t-h-c:t-h}$, not necessarily all of $s_{t-h-d+1:t-h}$. So, some of the value of $\alpha_{f}$ can be informally called the ”advantage” $\pi_{h}$ gets over $\pi_{h+d}$, but not necessarily all of it. Nevertheless, our proofs will be using this formal definition.
Definition (Maximum Inference Disadvantage ($\epsilon_{b}$)). Consider the maximum divergence that can be accumulated by the ${(c,h)}$-model from not knowing the past states $s_{t-h-d-c:t-h-c-1}$ and let that be $\epsilon_{b}$. More formally, we say that, for fixed $C$ from Proposition 1, any state in $\mathcal{S}^{-}$ and any $z_{t-k:t}$ and any $\hat{s}_{t-h-d-c:t-h-c-1}\neq{s}_{t-h-d-c:t-h-c-1}$:

### v3

@@ 5.1 One-Dimensional Diagnostic Experiments
Result. As shown in Fig. 5, when the environment is deterministic ($\delta=0.0$), larger action horizons capture the expert distribution better, consistent with Corollary 2. With an action horizon of 10, the learner achieves zero total variation distance with the expert distribution. Conversely, when the environment is highly stochastic ($\delta=0.4$), an action horizon of 1 outperforms all other learners, corroborating with Corollary 3. With moderate noise ($\delta=0.2$), there is no monotonic pattern and the optimal policy turns out to be the one with action horizon 5. Refer to Sec. A.1 for more detailed results. This controlled experiment validates the tradeoff identified in Proposition 1, which we will further demonstrate in robotic manipulation subjected to stochastic noise in Sec. 5.2.3
@@ 5.2.1 Comparison with Existing Inference Methods
  • Exponential Moving Average (EMA) (Zhao et al., 2023b): Smooth action chunking by averaging a new prediction $a$ with the previous one $\hat{a}$ for each overlapping step $a_{t}=\lambda a_{t}+(1-\lambda){\hat{a}}_{t}$. This method is also known as temporal ensembling. By default, we set $\lambda=0.5$.
@@ 5.2.2 Scalability and Compatibility of BID
Stochastic Noise 0.0 1.0 1.5
Vanilla Open-Loop 64.0 26.9 13.0
BID Open-Loop 66.1 31.4 16.0
Vanilla Closed-Loop 48.9 38.3 29.5
BID Closed-Loop 54.4 45.3 31.7
Table 1: Success rates of VQ-BeT on the Push-T task under various noise conditions. Closed-loop BID substantially outperforms other methods in stochastic environments. See Table 4 for detailed ablations.
Sample Size Success (%) Time (ms)
1  (vanilla) 49.1 13.2
8  (ours) 52.9 25.2
16  (ours) 54.2 25.9
32  (ours) 54.4 26.8
@@ 6 Conclusion
Limitations. One major limitation of BID lies in its computational complexity. While the decoding computation can be fully parallelized on modern GPUs, it may remain expensive for high-frequency operations on low-cost robots. Designing algorithms that can generate quality yet diverse action chunks under batch size constraints can be an interesting avenue for future research. Additionally, our analysis and method have been limited to policies with short context lengths, driven by their empirical effectiveness with limited human demonstrations. Developing techniques capable of learning robust long-context policies can be another compelling direction for future research.
@@
Our code for the experiments with Diffusion Policy is available at https://github.com/YuejiangLIU/bid_diffusion and our code for the experiments with VQ-BET is available at https://github.com/Jubayer-Hamid/bid_lerobot. Additionally, detailed descriptions of our experimental setups are available in Appendix B, and complete proofs of our theoretical claims can be found in Appendix D.
@@ A.1 One-dimensional Simulations
Figure 11: Effect of action horizon ($h$) and context length ($c$) on diffusion policies in the Push-T task. The baseline is set at $h=2$ and $c=2$, with $\Delta h=h-c=0$. Extending the action horizon ($h>2$) consistently improves performance, whereas extending the context length ($c>2$) can cause substantial performance declines. Each model is trained for $5k$ epochs. Results are averaged over the last five checkpoints.
@@
Setup. Our work builds on the premise that the action horizon is longer than the context length, as commonly designed for recent policies. While BID mitigates the inherent limitations of this design choice through test-time decoding, an important question remains: could extending the context length yield stronger policies? To understand this, we trained diffusion policies with varying combinations of action horizons and context lengths on the Push-T task. Specifically, we use a short context length ($c=2$) and a short action horizon ($h=2$) as our baseline, set the prediction length equal to the action horizon $l=h$, and incrementally increase these parameters to larger values ${6,10,14}$ to assess their impact.
@@ A.3 Ablation Study of Forward Contrast
Figure 13: Effect of the decay rate for the exponential moving average. In each task, we measure the relative performance among different decay rates. The optimal decay rate varies by task, leading to a practical challenge of identifying a universal temporal ensembling strategy (Zhao et al., 2023b).
@@ A.4 Challenges for Temporal Ensebmling
A.5 Ablation Studies on VQ-BeT
Table 4 presents the performance of various sampling methods with VQ-BeT (Lee et al., 2024) on the Push-T task under different levels of environment stochasticity. We compare BID against vanilla sampling and EMA, for both open-loop and closed-loop executions. BID consistently outperforms the baselines, particularly in highly stochastic settings, with close-loop BID attaining the best performance on average. Moreover, the performance of EMA degrades significantly as the environment gets more noisy, likely due to the divergence in strategies between consecutive action chunks. Additionally, we conduct an ablation study on the forward contrast objective, isolating the roles of positive samples (from a strong model) and negative samples (from a weaker model). We observe that both types of samples contribute to the strong performance of BID across varying levels of environment stochasticity.
Stochastic Noise 0.0 1.0 1.5 Average
Vanilla 64.0 $\pm$ 4.2 26.9 $\pm$ 2.8 13.0 $\pm$ 0.4 34.6 $\pm$ 1.7
EMA 64.1 $\pm$ 1.7 27.6 $\pm$ 3.3 12.9 $\pm$ 1.1 34.9 $\pm$ 1.3
Backward (ours) 64.0 $\pm$ 1.3 27.6 $\pm$ 1.0 13.4 $\pm$ 2.4 35.1 $\pm$ 1.0
+ Positive (ours) 65.6 $\pm$ 1.9 29.7 $\pm$ 0.4 15.7 $\pm$ 1.7 37.0 $\pm$ 0.9
+ Negative (ours) 65.1 $\pm$ 2.7 26.6 $\pm$ 0.8 14.8 $\pm$ 2.4 35.5 $\pm$ 1.2
BID (ours) 66.1 $\pm$ 3.5 31.4 $\pm$ 3.0 16.0 $\pm$ 1.2 37.8 $\pm$ 1.6
((a)) Open-Loop Operation
Stochastic Noise 0.0 1.0 1.5 Average
Vanilla 48.9 $\pm$ 2.7 38.3 $\pm$ 3.4 29.5 $\pm$ 0.9 38.9 $\pm$ 1.5
EMA 52.6 $\pm$ 2.9 35.7 $\pm$ 2.2 18.4 $\pm$ 2.3 35.6 $\pm$ 1.4
Backward (ours) 52.7 $\pm$ 1.3 42.0 $\pm$ 3.2 29.4 $\pm$ 0.5 41.4 $\pm$ 1.2
+ Positive (ours) 53.8 $\pm$ 3.3 44.1 $\pm$ 2.5 30.4 $\pm$ 0.8 42.8 $\pm$ 1.4
+ Negative (ours) 53.0 $\pm$ 2.1 44.3 $\pm$ 2.2 30.8 $\pm$ 0.4 42.7 $\pm$ 1.0
BID (ours) 54.4 $\pm$ 1.8 45.3 $\pm$ 3.8 31.7 $\pm$ 0.3 43.8 $\pm$ 1.4
((b)) Closed-Loop Operation
Table 4: Success rates of VQ-BeT on the Push-T task under various conditions and sampling methods. The left table shows open-loop success rates, while the right table shows closed-loop success rates. BID consistently outperforms vanilla counterparts in both settings.
A.6 Choice of Distance Metrics
As described in Sec. 4.2, our method uses the L2 distance as the default distance measure for quantifying similarity between action chunks. Yet, L2 distance is just one of several viable options. To assess the sensitivity of our approach to the choice of distance metric, we extend the experiments in Sec. 5.2.1 to further evaluate BID with alternative choices, including L1 and cosine distance. Results from three robot manipulation tasks, summarized in Table 5, demonstrate that BID achieves substantial performance gains regardless of the metric used. In particular, BID with L1 distance surpasses the default L2 distance in performance, highlighting the untapped potential of our method.
Vanilla L2 (Ours) L1 (Ours) Cosine (Ours)
Square 0.68 $\pm$ 0.06 0.76 $\pm$ 0.03 0.76 $\pm$ 0.04 0.73 $\pm$ 0.04
Lift 0.12 $\pm$ 0.02 0.58 $\pm$ 0.06 0.68 $\pm$ 0.05 0.70 $\pm$ 0.02
Kitchen 0.22 $\pm$ 0.04 0.64 $\pm$ 0.05 0.70 $\pm$ 0.04 0.61 $\pm$ 0.03
Table 5: Effect of distance metric. We evaluate BID with different distance metrics on three robot manipulation tasks. Regardless of the chosen metric, BID substantially outperforms the vanilla baseline. Notably, the L1 distance yields an even stronger performance than the default L2 distance used in our other experiments.
@@ B.1.1 Environment Details
Our implementation of BID for VQ-BeT (Lee et al., 2024) is built upon the code of LeRobot (Cadene et al., 2024). We use the best public checkpoint as the strong policy and a checkpoint trained for 100k iterations as the weak policy. Since BID requires sample diversity, we set the temperature to be 0.5 for all methods in our experiments.
@@ B.2.1 Dynamic Placing
Following previous works (Chi et al., 2023; Prasad et al., 2024), we use a Franka Panda as the robot hardware and the vision-based diffusion policy for its operation. The robot is equipped with two cameras: one ego-centric camera mounted at the wrist of the robot, one third-person camera mounted at a static bracket. Both cameras provide visual observations at a resolution of $256\times 256$ pixels. The robot operates at a frequency of 10 Hz, with a prediction length of 16 time steps.
@@ Results.
Appendix C Additional Discussions
Relation to Option Discovery.
Action chunking and option discovery share similarities in modeling temporally extended actions. However, their designs and outcomes are often different. Option discovery typically aims to learn hierarchical policies, explicitly discovering high-level skills from low-level action sequences (Bagaria & Konidaris, 2020; Jiang et al., 2022; 2023; Zhao et al., 2023a). In contrast, action chunking operates directly on low-level action sequences with fixed horizons, bypassing the need for abstraction into high-level skills. This simplicity has proven effective in large-scale robotic foundation models (Team et al., 2024; Black et al., 2024), addressing the scalability challenge that option discovery has yet to overcome. Future work could explore variable-horizon option discovery as a compelling alternative to action chunking, potentially combining the benefits of temporal abstraction with the scalability demonstrated by action chunking.
Relation to Long Context.
Our analysis in Sec. 3 highlights the benefits of extending the action horizon to better capture temporal dependencies across actions. Another natural approach to capturing these dependencies is to extend the context length. However, long-context policies often suffer from robustness issues, due to spurious correlations between past and future actions, as evidenced in Appendix A2 and studied in de Haan et al. (2019); Wen et al. (2020). Nevertheless, the increasing availability of large-scale robotic datasets for pre-training and fine-tuning may help mitigate these challenges. Advances in leveraging long context may open up new opportunities in policy design, such as long-context transformers (Su et al., 2024) and recurrent neural networks (Hochreiter & Schmidhuber, 1997; Zhuo et al., 2020).
Appendix D Proofs
First, we establish the following lemma which will help us compare different function classes based on the information they have access to:
@@ Appendix D Proofs
where we used $c^{*}(X):=\arg\min_{c}\mathbb{E}_{X}[\mathcal{L}(f(X,Y),c)|X]$. We get the inequality by recognizing that $\mathbb{R}\subsetneq G$. For the reverse inequality, consider any $g(X)\in G$:
@@
D.1 Definitions
We, first, analyze the effect of reducing context horizon. We show that, provided action horizon is constant, decreasing context horizon causes performance of the optimal policy to decrease.
Consider a $(c,h)$-policy (i.e., the policy has context length $c$ and action horizon $h$) whose probability of taking action $a_{t}$ at time $t$ in a chunk generated at $t$ is referred to as
@@ D.1 Definitions
In particular, this is an equality when the state $s_{t-c-1}$, conditioned on $C$, takes on only one value almost surely.
@@
We refer to the class of functions that accept $a_{t}$ and $s_{t-c-1:t}$ as inputs as $X_{2}$. Similarly, the class of functions that do not accept $a_{t}$ as inputs but accept $s_{t-c-1:t}$ as inputs is $X_{1}$. The function class that accepts only $s_{t-c:t}$ and not $s_{t-c-1}$ or $a_{t}$ as inputs are elements of $X_{0}$. Lastly, the function class that accepts $s_{t-c:t}$ and $a_{t}$ as inputs, but not $s_{t-c-1}$, are elements of $X_{-1}$.
$\min_{\pi_{(c+1,h)}\in X_{2}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+1,h)},\pi^{*}\Big{|}C\right]$
$=\mathbb{E}_{a_{t}}\left[\min_{\pi_{(c+1,h)}^{\prime}\in X_{1}}\mathbb{E}_{s_{t-c-1}}\left[\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c+1,h)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t-c-1},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4)
$=\mathbb{E}_{a_{t}}\left[\mathbb{E}_{s_{t-c-1}}\left[\min_{\pi_{(c,h)}^{\prime}\in X_{0}}\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c,h)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t-c-1},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4)
$\leq\mathbb{E}_{a_{t}}\left[\min_{\pi_{(c,h)}^{\prime}\in X_{0}}\mathbb{E}_{s_{t-c-1}}\left[\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c,h)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t-c-1},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Jensen’s inequality)
$=\min_{\pi_{(c,h)}\in X_{-1}}\mathbb{E}_{a_{t}}\left[\mathbb{E}_{s_{t-c-1}}\left[\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*})\Big{|}a_{t},s_{t-c-1},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4).
Use the law of total expectation to conclude. The equality conditions for Jensen’s inequality provides us the equality condition for this relationship too. ∎
@@
$\alpha_{b}\coloneqq\min_{\pi_{(c,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (8)
where $C=\{s_{t-h-c:t-h},a_{t-h:t-1}\}$ and $G=\{a_{t},s_{t-k:t-h-c-1},s_{t-h+1:t},z_{t-k:t}\}$. Clearly, $\alpha_{b}\geq 0$. In particular, $\alpha_{b}=0$ when $s_{t-h-d-c:t-h-c-1}$ can be deterministically inferred by $\pi_{(c,h)}$ or when the expert policy is independent of them.
Definition (Maximum Inference Disadvantage ($\epsilon_{f}$)). Consider the maximum divergence that can be accumulated by the ${(c,h+d)}$-policy from not knowing the recent states at time steps $s_{t-h-d+1:t-h}$ and let that be $\epsilon_{f}$. More formally, we say that, for fixed $C$ from Proposition 1, any states in $\mathcal{S}^{-}$, any $z_{t-k:t}$ and any $\hat{s}_{t-h-d+1:t-h}\neq{s}_{t-h-d+1:t-h}$:
@@
Proposition 7.
Let $\mathcal{L}$ be a non-linear, convex function. Let $c<k$. Let $G:=\{a_{t},s_{t-k:t-c-1},s_{t},z_{t-k:t}\}$ and $C:=\{s_{t-c:t-1},a_{t-1}\}$. Then,
@@
In particular, this is an equality when the state $s_{t}$, conditioned on $C$, takes on only one value almost surely.
@@
The proof is similar to that of Proposition 6. We refer to the class of functions that accept $a_{t}$ and $s_{t-c:t}$ as inputs as $X_{2}$. Similarly, the class of functions that do not accept $a_{t}$ as inputs but accept $s_{t-c:t}$ as inputs is $X_{1}$. The function class that accepts only $s_{t-c:t-1}$ and not $s_{t}$ or $a_{t}$ as inputs are elements of $X_{0}$. Lastly, the function class that accepts $s_{t-c:t-1}$ and $a_{t}$ as inputs, but not $s_{t}$, are elements of $X_{-1}$.
$\min_{\pi_{(c,h)}\in X_{2}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*}\Big{|}C\right]$
$=\mathbb{E}_{a_{t}}\left[\min_{\pi_{(c,h)}^{\prime}\in X_{1}}\mathbb{E}_{s_{t}}\left[\mathbb{E}_{s_{t-k:t-c-1},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c,h)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4)
$=\mathbb{E}_{a_{t}}\left[\mathbb{E}_{s_{t}}\left[\min_{\pi_{(c-1,h+1)}^{\prime}\in X_{0}}\mathbb{E}_{s_{t-k:t-c-2},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c-1,h+1)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4)
$\leq\mathbb{E}_{a_{t}}\left[\min_{\pi_{(c-1,h+1)}^{\prime}\in X_{0}}\mathbb{E}_{s_{t}}\left[\mathbb{E}_{s_{t-k:t-c-1},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c-1,h+1)}^{\prime},\pi^{*})\Big{|}a_{t},s_{t},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Jensen’s inequality)
$=\min_{\pi_{(c-1,h+1)}\in X_{-1}}\mathbb{E}_{a_{t}}\left[\mathbb{E}_{s_{t}}\left[\mathbb{E}_{s_{t-k:t-c-1},z_{t-k:t}}\left[\mathcal{L}(\pi_{(c-1,h+1)},\pi^{*})\Big{|}a_{t},s_{t},C\right]\Big{|}a_{t},C\right]\Big{|}C\right]$ (Lemma 4).
Use the law of total expectation to conclude. The equality condition can be seen from the equality condition of Jensen’s inequality. ∎
@@
$\alpha_{f}=\min_{\pi_{(c,h+d)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h+d)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (10)
where $C=\{s_{t-h-d-c:t-h-d},a_{t-h-d:t-1}\}$ and $G=\{a_{t},s_{t-k:t-h-d-c-1},s_{t-h-d+1:t},z_{t-k:t}\}$. Clearly, $\alpha_{f}\geq 0$. In particular, $\alpha_{f}=0$ when $\pi_{(c,h+d)}$ can infer $s_{t-h-d+1:t-h}$ perfectly i.e. when the environment is completely static with $P_{f}=1$. This makes sense–in the static environment, observing these states does not provide any advantage since the optimal $\pi_{(c,h+d)}$ can infer these states anyway using the actions taken at those time steps.
Definition (Maximum Inference Disadvantage ($\epsilon_{b}$)). Consider the maximum divergence that can be accumulated by the ${(c,h)}$-model from not knowing the past states $s_{t-h-d-c:t-h-c-1}$ and let that be $\epsilon_{b}$. More formally, we say that, for fixed $C$ from Proposition 1, any states in $\mathcal{S}^{-}$, any $z_{t-k:t}$ and any $\hat{s}_{t-h-d-c:t-h-c-1}\neq{s}_{t-h-d-c:t-h-c-1}$: